OpenStudy (anonymous):

I need help with the steps to solve this please: This table shows 4 observations of a pair of variables (x,y). The variables x and y are positively correlated with a correlation coefficient of 0.99. Find the slope, b, of the least squares regression line, y=ax+b, for these data: x 1 2 3 4 y 1 3 7 10

OpenStudy (anonymous):

send a picture of it because i dont understond it

OpenStudy (phi):

\[\left[\begin{matrix}1 & 1 \\ 2 &1\\ 3 & 1\\ 4 & 1\end{matrix}\right]\left(\begin{matrix}m \\ b\end{matrix}\right)=\left(\begin{matrix}1 \\ 3\\7\\10\end{matrix}\right)\]

OpenStudy (phi):

that is Ax= y denoting A' as A transpose x= inv(A'A) A' y